# Inverse statistics of active matter trajectories to distinguish interaction kernel anisotropy from emergent correlations

### v1

@@ 4 Effect of velocity alignment on emergent correlations
$f_{al}(r,\theta)=\eta_{0}F_{0}(r)+\eta_{1}\,\varepsilon F_{1}(r)\cos\theta+\eta_{2}F_{2}(r)\cos 2\theta+\cdots,$ (35)
where $F_{\ell}(r)$ are the Bessel transforms computed previously for the alignment-repulsion force, $f_{ar}$. That is, alignment is assumed to have the same radial and angular shape as the attraction-repulsion force up to per-harmonic scalings. The alignment drift in Equation (34) will now contribute to the Fokker-Planck equation,
@@
see Appendix E for details. Thus, alignment enters additively, with the same harmonic content as $f_{ar}$, and the Fourier-Bessel block structure of Section 3 is unchanged. Alignment changes only by having the first harmonic in $f_{al}$ contributing a diagonal shift to the dipole block, exactly as a dipolar kernel, $f_{\text{ar}}$ does in Section 28, while leaving the side matrix of Section 3.2 unaltered at leading order. Defining $\widehat{A}_{1}(k)$ as
$\widehat{A}_{1}(k)=2\pi\int_{0}^{\infty}r\,\big(\eta_{1}\varepsilon F_{1}(r)\big)J_{1}(kr)\,dr,$
@@
$\alpha_{al,1}(k)=\frac{\rho_{0}k^{2}}{2}\,\widehat{A}_{1}(k),\qquad\alpha_{ar,1}(k)=\frac{\rho_{0}k^{2}}{2}\,\varepsilon\,\widehat{F}_{1}(k),$ (38)
and $\beta=kv_{0}/2$ as before. The side branch eigenvalues remain the same, as no new terms enter. Two consequences follow immediately.
@@
First, the thresholds needed to induce a dipole instability shift, as any $\eta_{1}\neq 0$ simply replaces the dipole shift that occurs due to attraction-repulsion by $\alpha_{ar,1}\!+\alpha_{al,1}$ in Equation (37). A positive alignment term given by $\eta_{1}$ therefore raises $\lambda_{\mathrm{dip}}$ across 0 at lower $v_{0}$ or density, whereas a negative $\eta_{1}$ (i.e., anti-alignment) depresses it. Second, the parameter window in which dipolar instabilities are depressed can be either created or erased introducing the dipole shift. Because the side matrix, $M_{2}$, is unchanged, the inequality that guarantees this window is still
conform Equation (30) with $\alpha\mapsto\alpha_{ar,1}+\alpha_{al,1}$. Thus, anti-alignment in the front-back sector ($\eta_{1}<0$) widens the window even when $f_{ar}$ is isotropic, while pro-alignment narrows or destroys it. In this minimal extension, alignment acts as a harmonic- selective renormalisation of the dipolar instability with no change to the side matrix, so the emergent pattern hinges on the sign and magnitude of $\alpha_{al,1}(k)$ at the most unstable $k$ already identified in Section 3.
Figure 4: Dipole-suppression window under dipolar alignment. The isotropic growth rate, $\lambda^{(0)}r(k)$ (solid) and the side threshold (dashed-dotted) define a band of wavenumbers where the side branch is unstable while the dipole is suppressed by the combined shift $\alpha_{\mathrm{ar},1}(k)+\alpha_{\mathrm{al},1}(k)$ as shown by Equations (35)-(37). Plots for the same parameter set giving the disperesion relations as Figure 3.
As an illustration, Figures 3 and 4 show the dispersion relations for the same base parameters as the suppressed dipole regime in Figure 1 with a weak dipolar component in the interaction kernel ($\epsilon=-0.5$, $A_{1}=1.0$, $L_{1}=1.0$). At these parameters, the Morse attraction–repulsion kernel alone produces a marginally stable side branch and a weakly unstable dipole branch, while the quadrupole mode remains strongly damped. We introduce a separate dipolar alignment term, with amplitude $\eta_{1}$. This term selectively shifts the dipole eigenvalue without affecting the side or quadrupole branches to leading order. We therefore explore $\eta_{1}\in\{0,\,+0.3,\,-0.3\}$, corresponding respectively to: no alignment, alignment, and anti-alignment, while all other parameters remain fixed. For $\eta_{1}=0$, the dispersion relation $\lambda_{\mathrm{dip}}(k)$ exhibits a shallow positive maximum near $k\simeq 1.2$, indicating a weak dipolar instability. For $\eta_{1}>0$ (alignment), $\alpha_{\mathrm{al},1}(k)>0$ and the dipole curve shifts upward, slightly enhancing this mode. For $\eta_{1}<0$ (anti-alignment), $\alpha_{\mathrm{al},1}(k)<0$ and the dipole curve shifts downward while the side branch $\lambda_{\mathrm{side},+}(k)$ remains unchanged, thereby creating a distinct band of $k$ for which
@@
corresponding to the dipole-suppressed regime shown in Figure 4. Put together, this demonstrates that an alignment term, rather than the dipolar interaction itself, that can stabilise the homogeneous state by selectively damping the dipole branch while leaving higher modes largely unchanged.

### v2

@@ 4 Effect of velocity alignment on emergent correlations
$f_{al}(r,\theta)=\eta_{0}F_{0}(r)+\eta_{1}\,\varepsilon F_{1}(r)\cos\theta,$ (35)
i.e. so that the Bessel transforms computed for the alignment kernel will be proportional to those of the attraction-repulsion kernel, $f_{\text{ar}}$. The alignment drift in Equation (34) will now contribute to the Fokker-Planck equation,
@@
see Appendix E for details. Thus, alignment enters additively, with the same harmonic content as $f_{ar}$, and the Fourier-Bessel block structure of Section 3 is unchanged. Alignment changes only by having the first harmonic in $f_{al}$ contributing a diagonal shift to the dipole block, exactly as a dipolar kernel, $f_{\text{ar}}$ does in Equation (28), while leaving the side matrix of Section 3.2 unaltered at leading order. Defining $\widehat{A}_{1}(k)$ as
$\widehat{A}_{1}(k)=2\pi\int_{0}^{\infty}r\,\big(F_{1}(r)\big)J_{0}(kr)\,dr,$
@@
$\alpha_{al,1}(k)=\frac{\rho_{0}k^{2}}{2}\eta_{1}\varepsilon\,\widehat{A}_{1}(k),\qquad\alpha_{ar,1}(k)=\frac{\rho_{0}k^{2}}{2}\,\varepsilon\,\widehat{F}_{1}(k),$ (38)
and $\beta=kv_{0}/2$ as before. The side branch eigenvalues remain the same, as no new terms enter. Two consequences follow immediately. First, the thresholds needed to induce a dipole instability shift, as any $\eta_{1}\neq 0$ simply replaces the dipole shift that occurs due to attraction-repulsion by $\alpha_{ar,1}\!+\alpha_{al,1}$ in Equation (37). A positive alignment term given by $\eta_{1}$ therefore raises $\lambda_{\mathrm{dip}}$ across 0 at lower $v_{0}$ or density, whereas a negative $\eta_{1}$ (i.e., anti-alignment) depresses it. Second, the parameter window in which dipolar instabilities are depressed can be either created or erased. Because the side matrix, $M_{2}$, is unchanged, the inequality that guarantees this window is still
cf. Equation (30) with $\alpha\mapsto\alpha_{ar,1}+\alpha_{al,1}$. Thus, anti-alignment in the front-back sector ($\eta_{1}<0$) widens this window even when $f_{ar}$ is isotropic, while pro-alignment narrows or destroys it. In this minimal extension, alignment can selectively modulate the dipole instability with no change to the side matrix, and the final resulting pattern hinges on the sign and magnitude of $\alpha_{al,1}(k)$.
@@
As an illustration, Figure 3 shows the dispersion relation for the same parameters as the suppressed dipole regime in Figure 1 with a weak dipolar component in the attraction-repulsion kernel ($\epsilon=-0.5$, $A_{1}=1.0$, $L_{1}=1.0$). At these parameters, the Morse attraction–repulsion kernel alone produced an unstable quadrupole instability as well as an unstable dipole instability (with a smaller growth rate). We now introduce a separate dipolar alignment term, with amplitude $\eta_{1}$. As discussed previously, this term can selectively shift the dipole eigenvalue without affecting the quadrupole growth rate. We therefore explore $\eta_{1}\in\{0,\,+0.3,\,-0.3\}$, corresponding respectively to: no alignment, alignment, and anti-alignment, while all other parameters remain fixed. For $\eta_{1}=0$, the dispersion relation $\lambda_{\mathrm{dip}}(k)$ exhibits a maximum of zero near $k\simeq 2.1$, indicating a critical point for an emergent dipole instability, as before. For $\eta_{1}>0$ (alignment), $\alpha_{\mathrm{al},1}(k)>0$ and the dipole curve shifts upward, slightly enhancing this mode. For $\eta_{1}<0$ (anti-alignment), $\alpha_{\mathrm{al},1}(k)<0$ and the dipole curve shifts downward while the side branch $\lambda_{\mathrm{side},+}(k)$ remains unchanged, thereby creating a distinct band of $k$ for which
@@
corresponding to a parameter regime in which the dipole instability is even further suppressed. Put together, this demonstrates that an alignment term can selectively dampen or strengthen the dipole instability while leaving higher modes largely unchanged, thus fundamentally changing the linear stability properties of the emergent correlation field.